# Trivial Isomorphisms between Reduced Products

### v5

@@ 2.2.2. Graphs
The canonical models for randomness introduced by Gilbert ([21], but see a random reference on random graphs such as [18]) are sufficiently random. To wit, if $n\geq 1$ and $10<p<$ 1, then $\mathbb{G}_{n,p}$ is a probability space of objects with $n$ vertices, and associated to the ($n\choose 2$ pairs of these vertices we have independent random variables $x_{ij}$, $1\leq i<j\leq n$, such that $x_{ij}=1$ with probability $p$. Given $x_{ij}$, vertices $i$ and $j$ are connected if and only if $x_{ij}=1$. This describes a probability space on the space of all simple graphs with $n$ vertices, and as $n\to\infty$ the probability that a graph in this space is sufficiently random converges to 1. The following is a consequence of standard facts about random graphs and it will play a role in the proof of Theorem 5.
@@ 2.3. A general local criterion
Let $T$ be a first order theory, and suppose that ${\mathcal{Z}}$ is a set of pairs definable in $T^{*}$. We say that $T$ recognizes coordinates at ${\mathcal{Z}}$ if for all $(a,b)\in{\mathcal{Z}}$ we have that $a\perp b$ and the sets $A_{ab}$ (as in (2.6)), as well as their canonical Boolean algebra structure, are uniformly definable131313‘Uniformly definable’ means that a single formula with parameters $a$ and $b$ defines $A_{ab}$ for all relevant pairs $a,b$ in all relevant reduced products, and that analogous fact applies to Boolean operations in the canonical Boolean algebra structure on $A_{ab}$. in all reduced products of models of $T$ over atomless ideals.
@@ 3. Open Colouring Axioms
In [11], the second author defined an apparent strengthening of $\operatorname{{\mathsf{OCA}_{T}}}$ named $\operatorname{{\mathsf{OCA}_{\infty}}}$. [^14] In [8] a weaker axiom was called $\operatorname{{\mathsf{OCA}_{\infty}}}$, although a proof that $\operatorname{{\mathsf{OCA}_{\infty}}}$ as introduced in [11] is a consequence of $\operatorname{{\mathsf{PFA}}}$ was given. (We should note that this proof contains a glaringly obvious typo; in 4. of Lemma 3.1, ‘whenever $s\subseteq t$’ should be ‘for some $t$ such that $s\subseteq t$.’) This axiom was proved to be equivalent to $\operatorname{{\mathsf{OCA}_{T}}}$ ([31, §5], see also [13, Theorem 8.6.6]). $\operatorname{{\mathsf{OCA}^{\#}}}$ can be viewed as a refined version of $\operatorname{{\mathsf{OCA}_{\infty}}}$, and Moore’s proof that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}_{\infty}}}$ easily adapts to prove that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}^{\#}}}$.
@@
To make the relation between $\operatorname{{\mathsf{OCA}^{\#}}}$ and $\operatorname{{\mathsf{OCA}_{T}}}$ (in the form of $\operatorname{{\mathsf{OCA}_{\infty}}}$) more obvious, the reader may want to take a look at the reformulation of $\operatorname{{\mathsf{OCA}_{T}}}$ in terms of uniformization of coherent families of functions given in [9, Proposition 2.2.11]. With $K_{0}^{n}=\bigcup{\mathcal{V}}_{n}$, the first alternatives of $\operatorname{{\mathsf{OCA}_{\infty}}}$ and $\operatorname{{\mathsf{OCA}^{\#}}}$ are equivalent, and the second alternative of the latter is a sharper variant of the second alternative of the former in which one keeps track of the witness (typically a basic open set), provided by $\rho$, for the reason why $\{f(a),f(b)\}$ belongs to $K^{n}_{0}$.
@@
Our proof of the following is based on [31, §5], and we are grateful to Boban Veličković for pointing out that $\operatorname{{\mathsf{OCA}^{\#}}}$ is probably a consequence of $\operatorname{{\mathsf{OCA}_{T}}}$.
@@ 3.1. A uniformization principle for families of partial selectors
In our applications all ${\mathcal{Y}}_{n}$ will be separable and metrizable (second countable), hence the topology on ${\mathcal{F}}$ will also be separable and metrizable (second countable). If all spaces ${\mathcal{Y}}_{n}$ are Polish then so is the space of partial selectors.
@@
For $s\in\{0,1\}^{<{\mathbb{N}}}$ let $|s|$ denote its length (equivalently, its domain).
@@
Assume $\operatorname{{\mathsf{OCA}_{T}}}$. If ${\mathcal{F}}$ is a family of partial selectors for a family ${\mathcal{Y}}_{n}$, for $n\in{\mathbb{N}}$, of second-countable topological spaces, then one of the following possibilities holds.
@@
(2): Enumerate the $n$th level of $\{0,1\}^{<{\mathbb{N}}}$ as $s(n,i)$, for $i<2^{n}$. Define the sets $S(n,i)$, for $n\in{\mathbb{N}}$ and $i<2^{n}$ by setting $S(n,i)=A(s(n,i))$ when $s(n,i)\in D$ and by choosing $S(n,i)$ to be an arbitrary set of cardinality $n+(4^{n+1}-1)/3$ when $s(n,i)\notin D$. By part (1) of this lemma, there are pairwise disjoint $F(n,i)\subseteq S(n,i)\setminus n$ for $n\in{\mathbb{N}}$ and $i<2^{n}$, of cardinality $2^{n}$. Then by defining for every $s\in D$ the set $B(s)=F(n,i)$, where $n$ and $i<2^{n}$ are such that $s=s(n,i)$, we find that the sets $B(s)\subseteq A(s)\setminus|s|$ are as required.
@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
It remains to prove that ${\mathcal{J}}_{\textrm{cont}}$ is nonmeager. Since the constant map sending $\mathcal{P}({\mathbb{N}})$ to the empty set is continuous and lifts $\Phi$ on $\operatorname{Fin}$, ${\mathcal{J}}_{\textrm{cont}}$ contains $\operatorname{Fin}$. We work by contradiction, and assume that ${\mathcal{J}}_{\textrm{cont}}$ is meager. By the well-known characterization of nonmeager ideals of Talagrand and Jalali-Naini ([9, Theorem 3.10.1]), there are pairwise disjoint finite intervals of ${\mathbb{N}}$, $F_{j}$, for $j\in{\mathbb{N}}$, such that $\bigcup_{j\in A}F_{j}\in{\mathcal{J}}_{\textrm{cont}}$ if and only if $A$ is finite. Fix any perfect tree-like almost disjoint family ${\mathcal{A}}_{0}$ and let
@@
Fix a lifting $\Phi_{*}$ of $\Phi$. A continuous homomorphism lifting an endomorphism of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ on some set $A$ is completely additive (see [9, Corollary 3.8.2]). This means that for each $A$ in ${\mathcal{J}}_{\textrm{cont}}$ the lifting of the restriction of $\Phi$ to ${\mathcal{P}}(A)/\operatorname{Fin}$ is of the form $B\mapsto h_{A}^{-1}(B)$ for a function
@@
A function between Polish spaces is called C-measurable if it is measurable with respect to the $\sigma$-algebra generated by analytic sets [^16] Kechris in [27, §29.D] calls such functions $\sigma(\Sigma^{1}_{1})$-measurable, while C-measurability denotes a weaker condition. Nevertheless, we stick to this terminology as it is commonly used in the area of study of liftings (e.g., [13, §17], [30], or [12]).. In the following proof, we use the Jankov–von Neumann uniformization theorem, stated for convenience in Appendix A.
@@
Since $\mathcal{W}$ is nonmeager and analytic, there is a basic open subset $U$ of ${\mathcal{P}}({\mathbb{N}})$ such that $U\setminus\mathcal{W}$ is meager (see e.g. [27, Proposition 8.26]). It is of the form $U=\{A\mid A\cap J_{0}=s_{0}\}$ for some finite interval $J_{0}$ and $s_{0}\subseteq J_{0}$. As $\mathcal{W}$ is hereditary, we may assume that $s_{0}=\emptyset$.
@@
We claim that $\operatorname{Diff}(\Theta(A),g_{A})$ is finite for every $A\in{\mathcal{J}}_{\textrm{cont}}$. Assume otherwise and fix $A\in{\mathcal{J}}_{\textrm{cont}}$ for which this set is infinite. Since ${\mathcal{X}}_{n}$ is nonmeager, its hereditary closure $\hat{\mathcal{X}}_{n}=\{B\subseteq{\mathbb{N}}\mid(\exists C\in{\mathcal{X}}_{n})B\subseteq C\}$ is nonmeager and hereditary and by [9, Theorem 3.10.1] there is an infinite $B\in\hat{\mathcal{X}}_{n}$ included in $\operatorname{Diff}(\Theta(A),g_{A})$. If $C\in{\mathcal{X}}_{n}$ with $B\subseteq C$, then $B\subseteq\operatorname{Diff}(g_{A},g_{C})\cup\operatorname{Diff}(g_{C},\Theta(A))$. However, each one of these two sets is finite: $\operatorname{Diff}(g_{A},g_{C})$ is finite because both $A$ and $C$ belong to ${\mathcal{J}}_{\textrm{cont}}$ and $|\operatorname{Diff}(g_{C},\Theta(A))|\leq 3n$ by Claim 4.8; contradiction.
@@
is a C-measurable lifting of $\Phi$ on ${\mathcal{J}}_{\textrm{cont}}$. By [9, Lemma 3.3.4], $\Phi$ decomposes as $\Phi_{1}\oplus\Phi_{2}$ where $\Phi_{1}$ has a continuous lifting and the kernel of $\Phi_{2}$ is nonmeager. By [9, Theorem 1.6.1], $\Phi_{1}$ has a completely additive lifting, as required. ∎
The conclusion of Theorem 4.7 is incompatible with CH. An obvious route towards strengthening it is by requiring $\Phi$ to have a continuous lifting; this strengthening was even named Strong Extension Principle and conjectured to follow from forcing axioms in [9, §4] (the Weak Extension Principle is a generalization of the conclusion of Theorem 4.7—or rather, its dual restatement in terms of Boolean algebras, see [13, §1.3.1]—to Čech–Stone remainders of locally compact Polish spaces other than ${\mathbb{N}}$). This conjecture is, however, false. By [6], there is an endomorphism of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ which does not have a lifting which is an endomorphism of ${\mathcal{P}}({\mathbb{N}})$.
@@ 5. Proof of Theorem 7
Each of the three subsections of the present section is centered around a proposition. These three, altogether, will give the main result of this section:
@@
is a coordinate respecting function. Then $\Phi$ is of twisted product form.
@@
Let $\Phi\colon\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}\to\prod_{n}{\mathcal{N}}_{n}/\operatorname{Fin}$ be a coordinate respecting function. By definition, we have an isomorphism $\alpha=\alpha^{\Phi}\colon{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}\to{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ such that for every $X\in{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ there exists a function $\Phi_{X}\colon{\mathcal{M}}\restriction X\to{\mathcal{N}}\restriction\alpha(X)$ that makes the following diagram commute
@@
is nonmeager, thus proving a local version of Theorem 5.1. In §5.2 we will use $\operatorname{{\mathsf{OCA}_{T}}}$ to uniformize local liftings associated with elements of ${\mathcal{J}}_{\rm{prod}}$ and obtain a Borel-measurable lifting of $\Phi$. Finally, in Proposition 5.10 we will prove that if $\Phi$ has a Baire-measurable lifting then it is of product form.
@@
is a coordinate respecting function with $\alpha^{\Phi}=\operatorname{id}$, then the ideal ${\mathcal{J}}_{\rm{prod}}$ is nonmeager.
@@
We will prove that for every tree-like almost disjoint family ${\mathcal{A}}$ all but countably many of its elements belong to ${\mathcal{J}}_{\rm{prod}}$ (see (5.1)). Using the well-known characterization of nonmeagerness of ideals on ${\mathbb{N}}$ due to Jalali-Naini and Talagrand ([9, Theorem 3.10.1]) again, this will imply that the ideal ${\mathcal{J}}_{\rm{prod}}$ is nonmeager.
@@
Let $\{(a_{\xi},A_{\xi}):\xi<\omega_{1}\}$ be a one-to-one enumeration of $H$ (by having passed to a subset if necessary, we may assume that $H$ has cardinality $\aleph_{1}$). Consider the sets $E_{\xi}=\{q\in\mathbb{P}\mid(\exists\eta>\xi)(a_{\eta},A_{\eta})\in F_{q}\}$, for $\xi<\omega_{1}$, and $E_{n}=\{q\in\mathbb{P}\mid n_{q}>n\}$, for $n\in{\mathbb{N}}$. Each $E_{n}$ is dense in $\mathbb{P}$, as can be seen from the proof of Claim 5.3. Each one of the $E_{\xi}$ is a nonempty subset of $\mathbb{P}$. Since $\mathbb{P}$ has the countable chain condition, some $p\in\mathbb{P}$ forces that the generic filter $\dot{G}$ intersects uncountably many of the $E_{\xi}$ nontrivially. (To see this, assume otherwise and fix a maximal antichain consisting of conditions that decide the supremum of the set of all $\xi$ such that $\dot{G}\cap E_{\xi}\neq\emptyset$. Since the antichain is countable we have $E_{\xi}=\emptyset$ for all but countably many $\xi$; contradiction.) Since $\xi<\eta$ implies $E_{\xi}\supseteq E_{\eta}$, each $E_{\xi}$ is dense below $p$.
@@
This proves that ${\mathcal{J}}_{\rm{prod}}$ intersects ${\mathcal{A}}$ nontrivially. Since every subfamily of $\mathcal{A}$ is tree-like, and therefore if it is in addition uncountable then ${\mathcal{J}}_{\rm{prod}}$ intersects it, ${\mathcal{J}}_{\rm{prod}}$ includes all but countably many elements of ${\mathcal{A}}$. ∎

### v6

@@ 2.2.2. Graphs
The canonical models for randomness introduced by Gilbert ([21], see also a general reference on random graphs such as e.g. [18]) are sufficiently random. To wit, if $n\geq 1$ and $10<p<$ 1, then $\mathbb{G}_{n,p}$ is a probability space of objects with $n$ vertices, and associated to the ($n\choose 2$ pairs of these vertices we have independent random variables $x_{ij}$, $1\leq i<j\leq n$, such that $x_{ij}=1$ with probability $p$. Given $x_{ij}$, vertices $i$ and $j$ are connected if and only if $x_{ij}=1$. This describes a probability space on the space of all simple graphs with $n$ vertices, and as $n\to\infty$ the probability that a graph in this space is sufficiently random converges to 1. The following is a consequence of standard facts about random graphs and it will play a role in the proof of Theorem 5.
@@ 2.3. A general local criterion
Let $T$ be a first order theory, and suppose that ${\mathcal{Z}}$ is a set of pairs definable in $T^{*}$. We say that $T$ recognizes coordinates at ${\mathcal{Z}}$ if for all $(a,b)\in{\mathcal{Z}}$ we have that $a\perp b$ and the sets $A_{ab}$ (as in (2.6)), as well as their canonical Boolean algebra structure, are uniformly definable141414‘Uniformly definable’ means that a single formula with parameters $a$ and $b$ defines $A_{ab}$ for all relevant pairs $a,b$ in all relevant reduced products, and that analogous fact applies to Boolean operations in the canonical Boolean algebra structure on $A_{ab}$. in all reduced products of models of $T$ over atomless ideals.
@@ 3. Open Colouring Axioms
In [12], the second author defined an apparent strengthening of $\operatorname{{\mathsf{OCA}_{T}}}$ named $\operatorname{{\mathsf{OCA}_{\infty}}}$. [^15] In [9] a weaker axiom was called $\operatorname{{\mathsf{OCA}_{\infty}}}$, although a proof that $\operatorname{{\mathsf{OCA}_{\infty}}}$ as introduced in [12] is a consequence of $\operatorname{{\mathsf{PFA}}}$ was given. (We should note that this proof contains a glaringly obvious typo; in 4. of Lemma 3.1, ‘whenever $s\subseteq t$’ should be ‘for some $t$ such that $s\subseteq t$.’) This axiom was proved to be equivalent to $\operatorname{{\mathsf{OCA}_{T}}}$ ([32, §5], see also [14, Theorem 8.6.6]). $\operatorname{{\mathsf{OCA}^{\#}}}$ can be viewed as a refined version of $\operatorname{{\mathsf{OCA}_{\infty}}}$, and Moore’s proof that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}_{\infty}}}$ easily adapts to prove that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}^{\#}}}$.
@@
To make the relation between $\operatorname{{\mathsf{OCA}^{\#}}}$ and $\operatorname{{\mathsf{OCA}_{T}}}$ (in the form of $\operatorname{{\mathsf{OCA}_{\infty}}}$) more obvious, the reader may want to take a look at the reformulation of $\operatorname{{\mathsf{OCA}_{T}}}$ in terms of uniformization of coherent families of functions given in [10, Proposition 2.2.11]. With $K_{0}^{n}=\bigcup{\mathcal{V}}_{n}$, the first alternatives of $\operatorname{{\mathsf{OCA}_{\infty}}}$ and $\operatorname{{\mathsf{OCA}^{\#}}}$ are equivalent, and the second alternative of the latter is a sharper variant of the second alternative of the former in which one keeps track of the witness (typically a basic open set), provided by $\rho$, for the reason why $\{f(a),f(b)\}$ belongs to $K^{n}_{0}$.
@@
Our proof of the following is based on [32, §5], and we are grateful to Boban Veličković for pointing out that $\operatorname{{\mathsf{OCA}^{\#}}}$ is probably a consequence of $\operatorname{{\mathsf{OCA}_{T}}}$.
@@ 3.1. A uniformization principle for families of partial selectors
In our applications all ${\mathcal{Y}}_{n}$ will be separable and metrizable, hence the topology on ${\mathcal{F}}$ will also be separable and metrizable. If all spaces ${\mathcal{Y}}_{n}$ are Polish then so is the space of partial selectors.
@@
For $s\in\{0,1\}^{<{\mathbb{N}}}$ let $|s|$ denote its length (equivalently, its domain). Case (2) of Proposition 3.7 below mirrors one of the main technical contributions in [10], Lemma 3.13.5, where MA had been used to produce an almost disjoint family ${\mathcal{A}}$ and, for every $A\in{\mathcal{A}}$, an uncountable set of functions that pairwise disagree on $A$. We obtain analogous conclusions using $\operatorname{{\mathsf{OCA}_{T}}}$ (under the guise of $\operatorname{{\mathsf{OCA}^{\#}}}$) only.
@@
Assume $\operatorname{{\mathsf{OCA}_{T}}}$. If ${\mathcal{F}}$ is a family of partial selectors for a family ${\mathcal{Y}}_{n}$, for $n\in{\mathbb{N}}$, of separable metrizable topological spaces, then one of the following possibilities holds.
@@
(2): Enumerate the $n$th level of $\{0,1\}^{<{\mathbb{N}}}$ as $s(n,i)$, for $i<2^{n}$. Define the sets $S(n,i)$, for $n\in{\mathbb{N}}$ and $i<2^{n}$ by setting $S(n,i)=A(s(n,i))$ when $s(n,i)\in D$ and by choosing $S(n,i)$ to be an arbitrary set of cardinality $n+(4^{n+1}-1)/3$ when $s(n,i)\notin D$. By part (1) of this lemma, there are pairwise disjoint $F(n,i)\subseteq S(n,i)\setminus n$ for $n\in{\mathbb{N}}$ and $i<2^{n}$, of cardinality $2^{n}$. Then by defining for every $s\in D$ the set $B(s)=F(n,i)$, where $n$ and $i<2^{n}$ are such that $s=s(n,i)$, we find that the sets $B(s)\subseteq A(s)\setminus|s|$ are as required. ∎
@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
It remains to prove that ${\mathcal{J}}_{\textrm{cont}}$ is nonmeager. Since the constant map sending $\mathcal{P}({\mathbb{N}})$ to the empty set is continuous and lifts $\Phi$ on $\operatorname{Fin}$, ${\mathcal{J}}_{\textrm{cont}}$ contains $\operatorname{Fin}$. We work by contradiction, and assume that ${\mathcal{J}}_{\textrm{cont}}$ is meager. By the well-known characterization of nonmeager ideals of Talagrand and Jalali-Naini ([10, Theorem 3.10.1]), there are pairwise disjoint finite intervals of ${\mathbb{N}}$, $F_{j}$, for $j\in{\mathbb{N}}$, such that $\bigcup_{j\in A}F_{j}\in{\mathcal{J}}_{\textrm{cont}}$ if and only if $A$ is finite. Fix any perfect tree-like almost disjoint family ${\mathcal{A}}_{0}$ and let
@@
Fix a lifting $\Phi_{*}$ of $\Phi$. A continuous homomorphism lifting an endomorphism of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ on some set $A$ is completely additive (see [10, Corollary 3.8.2]). This means that for each $A$ in ${\mathcal{J}}_{\textrm{cont}}$ the lifting of the restriction of $\Phi$ to ${\mathcal{P}}(A)/\operatorname{Fin}$ is of the form $B\mapsto h_{A}^{-1}(B)$ for a function
@@
A function between Polish spaces is called C-measurable if it is measurable with respect to the $\sigma$-algebra generated by analytic sets [^17] Kechris in [28, §29.D] calls such functions $\sigma(\Sigma^{1}_{1})$-measurable, while C-measurability denotes a weaker condition. Nevertheless, we stick to this terminology as it is commonly used in the area of study of liftings (e.g., [14, §17], [31], or [13]).. In the following proof, we use the Jankov–von Neumann uniformization theorem, stated for convenience in Appendix A.
@@
Since $\mathcal{W}$ is nonmeager and analytic, there is a basic open subset $U$ of ${\mathcal{P}}({\mathbb{N}})$ such that $U\setminus\mathcal{W}$ is meager (see e.g. [28, Proposition 8.26]). It is of the form $U=\{A\mid A\cap J_{0}=s_{0}\}$ for some finite interval $J_{0}$ and $s_{0}\subseteq J_{0}$. As $\mathcal{W}$ is hereditary, we may assume that $s_{0}=\emptyset$.
@@
We claim that $\operatorname{Diff}(\Theta(A),g_{A})$ is finite for every $A\in{\mathcal{J}}_{\textrm{cont}}$. Assume otherwise and fix $A\in{\mathcal{J}}_{\textrm{cont}}$ for which this set is infinite. Since ${\mathcal{X}}_{n}$ is nonmeager, its hereditary closure $\hat{\mathcal{X}}_{n}=\{B\subseteq{\mathbb{N}}\mid(\exists C\in{\mathcal{X}}_{n})B\subseteq C\}$ is nonmeager and hereditary and by [10, Theorem 3.10.1] there is an infinite $B\in\hat{\mathcal{X}}_{n}$ included in $\operatorname{Diff}(\Theta(A),g_{A})$. If $C\in{\mathcal{X}}_{n}$ with $B\subseteq C$, then $B\subseteq\operatorname{Diff}(g_{A},g_{C})\cup\operatorname{Diff}(g_{C},\Theta(A))$. However, each one of these two sets is finite: $\operatorname{Diff}(g_{A},g_{C})$ is finite because both $A$ and $C$ belong to ${\mathcal{J}}_{\textrm{cont}}$ and $|\operatorname{Diff}(g_{C},\Theta(A))|\leq 3n$ by Claim 4.8; contradiction.
@@
is a C-measurable lifting of $\Phi$ on ${\mathcal{J}}_{\textrm{cont}}$. By [10, Lemma 3.3.4], $\Phi$ decomposes as $\Phi_{1}\oplus\Phi_{2}$ where $\Phi_{1}$ has a continuous lifting and the kernel of $\Phi_{2}$ is nonmeager. By [10, Theorem 1.6.1], $\Phi_{1}$ has a completely additive lifting, as required. ∎
The conclusion of Theorem 4.7 is incompatible with CH. An obvious route towards strengthening it is by requiring $\Phi$ to have a continuous lifting; this strengthening was even named Strong Extension Principle and conjectured to follow from forcing axioms in [10, §4] (the weak Extension Principle is a generalization of the conclusion of Theorem 4.7—or rather, its dual restatement in terms of Boolean algebras, see [14, §1.3.1]—to Čech–Stone remainders of locally compact Polish spaces other than ${\mathbb{N}}$). This conjecture is, however, false. By [7], there is an endomorphism of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ which does not have a lifting which is an endomorphism of ${\mathcal{P}}({\mathbb{N}})$.
@@ 5. Proof of Theorem 7
Each of the two subsections of the present section is centered around a proposition. These two, together, will give the main result of this section:
@@
is a coordinate-respecting function. Then $\Phi$ is of twisted product form.
@@
Let $\Phi\colon\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}\to\prod_{n}{\mathcal{N}}_{n}/\operatorname{Fin}$ be a coordinate-respecting function. By definition, we have an isomorphism $\alpha=\alpha^{\Phi}\colon{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}\to{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ such that for every $X\in{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ there exists a function $\Phi_{X}\colon{\mathcal{M}}\restriction X\to{\mathcal{N}}\restriction\alpha(X)$ that makes the following diagram commute
@@
is nonmeager, thus proving a local version of Theorem 5.1. In §5.2 we will use $\operatorname{{\mathsf{OCA}_{T}}}$ to uniformize local liftings associated with elements of ${\mathcal{J}}_{\rm{prod}}$ and obtain a lifting of product form.
@@
is a coordinate-respecting function with $\alpha^{\Phi}=\operatorname{id}$, then the ideal ${\mathcal{J}}_{\rm{prod}}$ intersects every uncountable almost disjoint family. In particular ${\mathcal{J}}_{\rm{prod}}$ is nonmeager.
@@
We will prove that for every almost disjoint family ${\mathcal{A}}$ all but countably many of its elements belong to ${\mathcal{J}}_{\rm{prod}}$ (see (5.1)). Using the well-known characterization of nonmeagerness of ideals on ${\mathbb{N}}$ due to Jalali-Naini and Talagrand ([10, Theorem 3.10.1]) again, this will imply that the ideal ${\mathcal{J}}_{\rm{prod}}$ is nonmeager.
@@
Let $\{(a_{\xi},A_{\xi}):\xi<\omega_{1}\}$ be an enumeration of $H$ (by having passed to a subset if necessary, we may assume that $H$ has cardinality $\aleph_{1}$). Consider the sets $E_{\xi}=\{q\in\mathbb{P}\mid(\exists\eta>\xi)(a_{\eta},A_{\eta})\in F_{q}\}$, for $\xi<\omega_{1}$, and $E_{n}=\{q\in\mathbb{P}\mid n_{q}>n\}$, for $n\in{\mathbb{N}}$. Each $E_{n}$ is dense in $\mathbb{P}$, as can be seen from the proof of Claim 5.3. Each one of the $E_{\xi}$ is a nonempty subset of $\mathbb{P}$. Since $\mathbb{P}$ has the countable chain condition, some $p\in\mathbb{P}$ forces that the generic filter $\dot{G}$ intersects uncountably many of the $E_{\xi}$ nontrivially. (To see this, assume otherwise and fix a maximal antichain consisting of conditions that decide the supremum of the set of all $\xi$ such that $\dot{G}\cap E_{\xi}\neq\emptyset$. Since the antichain is countable we have $E_{\xi}=\emptyset$ for all but countably many $\xi$; contradiction.) Since $\xi<\eta$ implies $E_{\xi}\supseteq E_{\eta}$, each $E_{\xi}$ is dense below $p$.
@@
This proves that ${\mathcal{A}}\setminus{\mathcal{J}}_{\rm{prod}}$ is countable and in particular that ${\mathcal{J}}_{\rm{prod}}$ intersects ${\mathcal{A}}$ nontrivially. ∎